OpenStudy (futureboy1):

convert 6.78 to a fraction= i will give medals and a fan

OpenStudy (dogzcatz):

678/100 and 339/50

OpenStudy (futureboy1):

thanks

OpenStudy (futureboy1):

how did you get that answer @DogzCatz

OpenStudy (futureboy1):

@kweenlexis

OpenStudy (kweenlexis):

6.78 = 678/100 = 339/50 = 6 39/50

OpenStudy (futureboy1):

thanks but how did u get that answer

OpenStudy (futureboy1):

kween this is really the question

OpenStudy (futureboy1):

Which mixed number is greater than 6.78? A. B. C. D.

OpenStudy (futureboy1):

sorry

OpenStudy (kweenlexis):

Write the decimal number as a fraction (over 1) 6.78 = 6.78 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 6.78 × 10 × 10 = 678 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 678 / 100 Simplifying our fraction = 678/100 = 339/50 = 6 39/50

OpenStudy (kweenlexis):

that was the question?

OpenStudy (futureboy1):

a is 6 2/3

OpenStudy (futureboy1):

b is 6 3/4

OpenStudy (futureboy1):

c is 6 7/9

OpenStudy (futureboy1):

that is the question

OpenStudy (futureboy1):

d is 6 4/5

OpenStudy (kweenlexis):

give me a second im taking a test :)

OpenStudy (dogzcatz):

I don't know I learned it in school Pretty much it's a % so that number out of a hundred divide by a number to get your answer